Skip to Search
|Library Catalog | Databases | Site Map | Search|
|Information Technology Who We Are Meeting Mins Equip Checkout Report Probs FAQs Phone Ext|
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Jason Fleming, Winston Harris, Michael Jay, Paul McDonough, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Stephen WilliamsAnnouncements
Ongoing IT Updates
- Michael Jay had a question about helping the PPD guy, Jeff, with IT issues. He needs to contact the PPD IT department regarding such matters.
- Amy got her car back.
- Orders - nothing significant.
- Batteries and charger returned to Lenovo to determine if the issue is with the batteries or charger. They've sent them on to Datamation for evaluation.
Workstation Support Unit
- Tivoli testing - stumbling block - Allen Rout came over last week and tested the archiving component for DLC. There will be a cost component to upgrade the Linux file system. Allen is checking on the cost. The rest is ready to go - right now we are backing up the file server, Web and Greenstone to CNS.
- SAN migration - Logan received the documentation from Angela for the West migration. We will do it during the evening hours. Administration should be pretty straightforward. After that, we'll be left with the Projects folder.
- CJK machine is almost ready to go. Logan is testing it with the MS Steady State software, rather than Deep Freeze. Will give it back to HSS and see how they like it.
- MS Steady State testing - Windows disk protection can't unfreeze on computers. We would need a way to unfreeze multiple machines. The control part is easier and better. It will be a couple of weeks before decisions can be made.
- Will spoke about AD migration to Central Exchange at the Cross-Divisional meeting. We've got a meeting next Tuesday to get it rolling.
- Disaster Prep - need to work on automated shutdown. UPS initiated shutdown - can do a shutdown script for storage units.
- McAfee - tiered scanning - don't need on access scan, can set it to scan C: drive.
- Public workstation cleanup complete (post virus).
- We are waiting on an image for West. Should be ready by next week.
- Patron logging and loop back policy discussion.
- We need to speed up the public machines.
- WSUS is working on Campus AD. When we're on AD, we will still use McAfee.
Software Projects Status
- Jason made a presentation about ALEPH yesterday.
- Brittle request is now in the catalog.
- Logan has the VIVO server set up - and remote SQL server. He is getting VIVO files installed. We will test before it goes to Val and Jana. Logan is getting quotes for parts ready to be submitted.
- Grover mods - small - it now shows the name of the department that the Requester resides.
- ARL stats - Michael Jay - finished the documentation and will start the database design.
- Bryant - Chris finished EAD tool. Next step is to load into UFDC.
- UF is switching to the new financial system this week. Winston is getting a lot of requests for staff PeopleSoft roles for those that access student financials and FERPA data. Training is required and they need to have their FERPA certification attached in Grover. We need to make sure Grover attachments are backed up.
IT Security Management
- Tom fixed Christian's problem with Contribute.
- Emerging Technologies group - Jason Fleming, Matthew Daley, Tom Minton, Ellie from Law, Elizabeth Outler (HSC), Laurie Taylor, Laura Jordan, Valrie Davis and Missy Shoop. Michelle Crump is the Chair. They are looking at suggestions like Second Life, etc to see how they fit into the scope of the Libraries. The following groups have disbanded: Resource Navigation, Web Advisory and Web 2.0. They are very interested in the X server - FCLA can run it, results from MetaLib can be repackaged. It can be slick.
- VIVO (Gator Scholar) - met with Mike Conlon. He is very interested. He'd like to see Tenure & Promotion tied into it.
- Discussions about redoing the Library Web page. Templates not mandatory. Library West (Stacey Ewing, Leilani & Laura Jordan) - working on their site.
- Stephen is working on a customizable homepage.
- Steve Carrico reported that an old version of the Acquisitions Web site was up and he needed the more current version restored. Tom was able to do so.
- Logan - Blog server upgraded to new version. Games work better. Cleaned up the firewall, configured FTP, and created accounts for the Web programmers.
- Web site search engine for the Library page has been narrowed.
- Meeting held.